OpenStudy (anonymous):

If f(3) = 9 and f '(x) ≥ 2 for 3 ≤ x ≤ 7, how small can f(7) possibly be?

OpenStudy (freckles):

Hmm seems like a mean value problem

OpenStudy (freckles):

\[c \in (a,b) \\ f'(c)=\frac{f(b)-f(a)}{b-a} \\ f'(c)=\frac{f(b)-f(a)}{b-a} \ge 2 \] input your numbers and solve for f(7)

OpenStudy (freckles):

@nonhacker Hope that helps I have to go peace